OpenStudy (starryknight):

please help me :( Solve by factoring. 3x2 + 11x - 20 = 0 answer choices; x = 5 or x = -1.33333 x = -5 or x = 1.33333 x = -1.66667 or x = 4 x = -0.25 or x = -1.66667

OpenStudy (anonymous):

okay so you can write 3x^2+11x-20 as 3x^2+15x-4x-20=0 so 3x(x+5)-4(x+5)=0 so (x+5)(3x-4)=0 so x=-5 or x=4/3

OpenStudy (anonymous):

4/3 = 1.33333333333

OpenStudy (anonymous):

so x=-5 or x=1.3333333

OpenStudy (starryknight):

thnx :)

OpenStudy (anonymous):

welcome ^^